OpenStudy (anonymous):

How do you factor this completely x^12-8

OpenStudy (anonymous):

Is this multiple choice?

OpenStudy (anonymous):

no i have to explain or show work i tried but idk

OpenStudy (anonymous):

Well, the farthest I could get was\[x^{12} - 8 = (x^{4} - 2)(x^{8} + 2x^{4} + 4)\]

OpenStudy (anonymous):

That was done using difference in cubes

OpenStudy (turingtest):

Calcmathlete use the formula\[a^3-b^3=(a-b)(a^2+ab+b^2)\]probably by noticing that \(8=2^3\) is a perfect cube